2 consecutive even numbers whose product is 48
could u list the steps and answer?

Answer by mananth(16946)

let the numbers be x, x+2
...
x(x+2)=48
x^2+2x=48
-48
x^2+2x-48=0
x^2+8x-6x-48=0
x(x+8)-6(x+8)=0
(x+8)(x-6)=0
x= 6 OR -8
the numbers are 6,8
OR
-8,-6